OpenStudy (anonymous):

Two non-simultaneous events have an equal likelihood of occurring (50%). If the events are independent of each other, what is the probability that they will both occur

OpenStudy (anonymous):

For independent events \(A\) and \(B\), \[P(A\cap B)=P(A)P(B)=\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{4}\]